# The reason peculiar velocities grow faster in general relativity than in Newtonian gravity

### v1

@@ 1 Introduction
This work provides a direct comparison between the Newtonian/quasi-Newtonian and the relativistic analysis of linear peculiar velocities on a Friedmann-Robertson-Walker (FRW) background with flat spatial sections. We set the theoretical framework of the two approaches, identify the point where they start to diverge from each other and explain why this happens. We then derive the linear evolution laws of the associated peculiar-velocity fields, highlight the differences between the corresponding solutions and discuss their implications for addressing the ongoing bulk-flow question. In the process, we also highlight the role of the acceleration/4-acceleration as the driving force of linear peculiar velocities. More specifically, we show that the stronger the 4-acceleration, the faster the driven peculiar-velocity field. Finally, we demonstrate a way of recovering the relativistic results from the Newtonian setup, by selectively accounting for certain source-free terms in Poisson’s formula that are typically sidestepped. Although, this recovery is rather ad hoc and not uniquely determined it provides useful analogies between the two types of study. For instance, one could argue that ignoring the above mentioned source-free terms in the Newtonian studies, is analogous to bypassing the purely general relativistic gravitational input of the peculiar flux.
@@ 2 Newtonian treatment of peculiar velocities
Given that and also depend on time, after a relatively straightforward calculation, the temporal derivative of the above leads to the linear auxiliary relation
@@ 3.3.1 Newtonian treatment
Newtonian peculiar-velocity studies start from differential equation (17). Recalling that to zero order, the linearised time derivative of the latter leads to
@@ 3.3.6 On the impact of the late-time acceleration
So far our analysis has been confined to the Einstein-de Sitter epoch of the universe. Within the framework of the CDM scenario the earlier dust-dominated era is followed by a recent phase of accelerated expansion, where the energy density of the cosmos is typically dominated by a cosmological constant, or by dark energy. Therefore, our results apply to the period between matter–radiation equality and the onset of cosmic acceleration. Once the latter epoch has started, the evolution of the peculiar-velocity field is expected to change drastically. More specifically, the accelerated expansion is expected to suppress the growth of peculiar velocities, in the same way it suppresses essentially all types of perturbations, like those in the matter density for example. This claim is also supported by numerical studies showing that dark energy inhibits the growth of peculiar velocities compared to the Einstein-de Sitter epoch [25]. Moreover, complementary studies looking into the linear evolution of peculiar velocities during a phase of de Sitter inflation, showed exponential decay for the -field [18]. Nevertheless, some at least of the strong earlier growth of the Einstein-de Sitter period should survive through the accelerated phase to the present. On these grounds, one should expect to observe residual bulk flows faster and deeper than those anticipated by the Newtonian studies, but also to see a drop in their mean velocity at lower redshifts [22]. Intriguingly, such a late-time suppression of the -field has been reported in the surveys of [8, 9].

### v2

@@ 1 Introduction
This work provides a direct comparison between the Newtonian/quasi-Newtonian and the relativistic analysis of linear peculiar velocities on a Friedmann-Robertson-Walker (FRW) background with flat spatial sections. The linear nature of all the studies implies that their results formally apply to relatively large scales, typically near and beyond the 100 Mpc threshold, where one feels confident enough to ignore the nonlinear effects.
We begin by setting the theoretical framework of the approaches, identify the point where they start to diverge from each other and explain why this happens. We then derive the linear evolution laws of the associated peculiar-velocity fields, highlight the differences between the corresponding solutions and discuss their implications for addressing the ongoing bulk-flow question. In the process, we also highlight the role of the acceleration/4-acceleration as the driving force of linear peculiar velocities. More specifically, we show that the stronger the 4-acceleration, the faster the driven peculiar-velocity field. Finally, we demonstrate a way of recovering the relativistic results from the Newtonian setup, by selectively accounting for certain source-free terms in Poisson’s formula that are typically sidestepped. Although, this recovery is rather ad hoc and not uniquely determined it provides useful analogies between the two types of study. For instance, one could argue that ignoring the above mentioned source-free terms in the Newtonian studies, is analogous to bypassing the purely general relativistic gravitational input of the peculiar flux.
@@ 2 Newtonian treatment of peculiar velocities
Given that and also depend on time, after a relatively straightforward calculation, the temporal derivative of the above leads to the linear time-evolution equation for , namely
@@ 3.3.1 Newtonian treatment
Newtonian peculiar-velocity studies start from differential equation (17). Recalling that to zero order, the linearised time derivative of the former leads to
@@ 3.3.6 On the impact of the late-time acceleration
So far our analysis has been confined to the Einstein-de Sitter epoch of the universe. Within the framework of the CDM scenario the earlier dust-dominated era is followed by a recent phase of accelerated expansion, where the energy density of the cosmos is typically dominated by a cosmological constant, or by dark energy. Therefore, our results apply to the period between matter–radiation equality and the onset of cosmic acceleration. Once the latter epoch has started, the evolution of the peculiar-velocity field is expected to change drastically. More specifically, the accelerated expansion is expected to suppress the growth of peculiar velocities, in the same way it suppresses essentially all types of perturbations, like those in the matter density for example.
Analytical relativistic studies that directly address and quantify the impact of a late CDM phase on large-scale peculiar velocities would be very useful indeed. At present, there is support for the aforementioned suppressing effect of the universal acceleration, but indirect.
In particular, according to the Newtonian numerical work of [25], dark energy inhibits the growth of peculiar velocities, when compared to the Einstein-de Sitter epoch. Also, studies looking into the linear evolution of peculiar velocities during a phase of de Sitter inflation, showed exponential decay for the -field [18]. Nevertheless, even if peculiar velocities are suppressed during the late -dominated phase, some at least of the strong growth of the preceding Einstein-de Sitter period should survive through to the present.
On these grounds, one should expect to measure bulk velocities faster than anticipated by the Newtonian studies at higher redshifts, but also to see a decrease in their mean value at lower redshifts (see [22] for further discussion). Here, we note that a peculiar-velocity profile, with increased values at higher redshifts (i.e. at earlier times) and a late-time suppression has been reported in the survey of [6] and (more pronounced) in those of [8, 26]. Put another way, our work appears to provide theoretical support to the aforementioned observational studies.
@@ 5 Discussion
Observationally, the time evolution of the peculiar velocities can be mapped into a redshift evolution at low , enabling a comparison between our theoretical predictions and the standard Newtonian expectation using peculiar-velocity surveys such as CosmicFlows-4. However, peculiar-velocity measurements are subject to a well-known degeneracy between the assumed background cosmology and local kinematic contributions. This degeneracy can be partially mitigated through independent distance indicators, such as Type Ia supernovae or Tully-Fisher relations, which provide redshift-independent distance estimates. Nevertheless, the use of such indicators introduces additional uncertainties that grow with redshift, thereby impacting the reconstruction of the peculiar-velocity field and limiting the robustness of the physical inferences. A detailed implementation of unbiased and statistically robust estimators — such as the minimum-variance reconstruction techniques applied to CosmicFlows-4 [8, 26] — is beyond the scope of the present work and will be addressed in future studies.